# Health Text Simplification: An Annotated Corpus for Digestive Cancer Education and Novel Strategies for Reinforcement Learning

### v2

@@ Abstract
Objective: The reading level of health educational materials significantly influences the understandability and accessibility of the information, particularly for minoritized populations. Many patient educational resources surpass the reading level and complexity of widely accepted standards. There is a critical need for high-performing text simplification models in health information to enhance dissemination and literacy. This need is particularly acute in cancer education, where effective prevention and screening education can substantially reduce morbidity and mortality.
Methods: We introduce Simplified Digestive Cancer (SimpleDC), a parallel corpus of cancer education materials tailored for health text simplification research, comprising educational content the American Cancer Society, Centers for Disease Control and Prevention, and National Cancer Institute. Utilizing SimpleDC alongside the existing Med-EASi corpus, we explore Large Language Model (LLM)-based simplification methods, including fine-tuning, reinforcement learning (RL), reinforcement learning with human feedback (RLHF), domain adaptation, and prompt-based approaches. Our experimentation encompasses Llama 2 and GPT-4. A novel RLHF reward function is introduced, featuring a lightweight model adept at distinguishing between original and simplified texts, thereby enhancing the model’s effectiveness with unlabeled data.
Results: Fine-tuned Llama 2 models demonstrated high performance across various metrics. Our innovative RLHF reward function surpassed existing RL text simplification reward functions in effectiveness. The results underscore that RL/RLHF can augment fine-tuning, facilitating model training on unlabeled text and improving performance. Additionally, these methods effectively adapt out-of-domain text simplification models to targeted domains. The best-performing RL-enhanced Llama 2 model outperformed GPT-4 in both automatic metrics and manual evaluation by subject matter experts.
@@ 1 Introduction
Accessibility and comprehensibility of patient medical educational materials are critical to improving health literacy and patient health [1]. Patient educational materials, including cancer-related materials, are essential to provide patients with the information necessary to make informed healthcare decisions. Health literacy is strongly influenced by the clarity and understandability of such materials and plays a key role in health outcomes [2]. It affects various aspects of patient health, including reducing risk factors, encouraging timely screening, and facilitating proactive health management. To ensure comprehensibility for a broad audience, the American Medical Association (AMA) and the National Institutes of Health (NIH) recommend a sixth-grade reading level or lower for patient educational materials [3, 4]. Despite the critical role of these educational materials in patient health, a substantial discrepancy persists between the recommended readability levels and the actual complexity observed in practice [5]. Educational materials from prominent sources are often at a high school or college reading level [5]. This reading complexity poses a barrier, particularly for people with lower literacy or limited health knowledge, preventing patients from fully understanding and taking action on essential health information. The accessibility and comprehensibility of educational materials are particularly concerning for disseminating cancer information, where clear and understandable guidance impacts patient education and decision-making [6, 7]. Our focus on text-based cancer educational materials stems from their pivotal role in patient education and the pressing need to improve their accessibility to diverse populations.
@@
The convergence of Natural Language Processing (NLP) and the evolution of Large Language Models (LLMs) offers a promising avenue to reconcile the disparity in reading levels in existing medical education materials through automatic simplification, to improve public health literacy [8]. LLMs have demonstrated proficiency in reducing linguistic complexity and preserving medical precision and meaning [9]. Finding the optimum balance of simplification and information retention is challenging, especially when the stakes of miscommunication are high [10]. Developing and evaluating data-driven text simplification models requires parallel corpora of original and simplified text. Creating these parallel corpora is extremely labor-intensive and costly. In the health domain, these challenges are exacerbated by the need for domain-specific medical expertise. To address these challenges, robust methods are needed to create high-performing simplification models with minimal training data.
@@
This paper makes significant contributions to the simplification of health information and text simplification more broadly. We focus on patient educational materials related to digestive cancer, given the prevalence and mortality of these cancers [11, 12]. We introduce Simplified Digestive Cancer (SimpleDC), a novel text simplification corpus comprising patient educational materials from three prominent sources: National Cancer Institute (NCI), American Cancer Society (ACS), and Centers for Disease Control (CDC). SimpleDC is unique in its concentration on digestive cancers and includes parallel sentences (original and simplified versions), where simplifications were generated by a team of nurse oncologists and a nurse practitioner with medical and patient education backgrounds. We explore supervised fine-tuning (SFT) and reinforcement learning (RL) approaches using Llama 2 [13] and introduce a novel approach for reinforcement learning with human feedback (RLHF). We also explore prompt-based methods using GPT-4 [14, 15] and introduce a novel self-correction strategy that utilizes emergent reasoning capabilities. Our evaluation includes standard automatic metrics (e.g. SARI, BLEU, BERTScore, and FKGL) and a human comparison of LLM- and human-generated simplifications. Automatic and manual evaluations demonstrate that the proposed RLHF reward function outperforms existing text simplification reward functions, complements SFT, and facilitates domain adaptation. Our findings advance health text simplification and enable the automatic creation of more accessible health information, ultimately contributing to improved health literacy. We provide the new SimpleDC corpus, trained models, and code to the research community. [^1] A link will be provided upon acceptance.
@@
Problem Existing patient education materials are frequently presented at a high literacy level, reducing accessibility for individuals with lower literacy.
What is already known Existing health text simplification corpora are limited, and there are knowledge gaps for text simplification related to reinforcement learning (RL), RL with human feedback (RLHF), and large language models (LLMs).
What this paper adds We present a novel text simplification corpus, Simplified Digestive Cancer (SimpleDC), and introduce an innovative RLHF reward function for training LLM simplification models. The findings demonstrate the RLHF approach surpasses existing methods in performance, complements supervised fine-tuning, and supports domain-specific adaptation.
@@ 2 Related Work
Text simplification involves the transformation of texts into versions that are more easily understandable to improve accessibility to various audiences. It has been explored in many domains (news articles, literature, scientific articles, etc.) where complex ideas must be communicated clearly and concisely [16, 17, 18, 19, 20]. Text simplification has been explored in the health domain through several tasks, such as generating patient-friendly summaries of clinical notes [21, 22, 23], making biomedical literature accessible to laypeople, and simplifying educational materials for patients [24, 25, 8, 26, 27, 28]. In the general domain, text simplification can focus on improving readability or making stylistic modifications; however, simplification of health information requires rigorous attention to meaning preservation and medical accuracy, given the critical implications of miscommunication [8, 29]. This challenge is amplified by the diverse needs of patients, which encompasses varying levels of literacy, language background, and other factors that require a nuanced approach to achieve clarity and accuracy.
Text simplification can be subdivided into three distinct tasks: lexical simplification, sentence-level simplification, and paragraph/document-level simplification [30, 31, 32]. These tasks address distinct challenges inherent in medical texts, such as medical jargon, complex sentence structures, and lengthy explanations. Lexical simplification commonly involves the substitution of words or phrases, including medical terminology [33, 34, 35, 29, 36]; however, the focus on word or phrase-level edits limits its scope and cannot accommodate more comprehensive rephrasing or restructuring.
@@
Sentence- and document-level simplification enables more comprehensive text restructuring to enhance overall clarity and accessibility and typically requires parallel text corpora (complex and simplified text) for model training and evaluation. Due to the labor-intensive nature of annotation, there are limited health-focused text simplification corpora, which we highlight here [37, 38, 9, 26]. The Medical dataset for Elaborative and Abstractive Simplification (Med-EASi) dataset comprises medical text from SIMPWIKI that was simplified by a team of medical experts and layperson crowd workers at the sentence-level [37]. The Autocomplete for Medical Text Simplification (AutoMeTS) dataset consists of automatically aligned sentences from English Wikipedia and Simple English Wikipedia for medical topics [26]. Devaraj et al. provided a manually annotated parallel corpus of biomedical literature with simplified text created by domain experts [9]. We introduce a new health text simplification corpus, SimpleDC, which is based on educational patient materials from prominent information sources (ACS, CDC, and NCI) and explicitly focuses on the cancer domain.
@@
Sentence-level simplification aims to enhance clarity by restructuring and rephrasing individual sentences. Early work used heuristics and traditional machine learning techniques, like pruning or pattern-based paraphrasing [39, 40]. Contemporary approaches conceptualize text simplification similar to machine translation, employing neural language models to transform sentences into simpler counterparts [8, 41]. Similarly, document-level simplification adopts a machine translation perspective [42, 43]. Essentially, all state-of-the-art text simplification models [44, 45] are based on LLMs, including: i) encoder-decoder architectures, like Text-to-Text Transfer Transformer (T5) [46] and Fine-tuned LAnguage Net on T5 (FlanT5) [47] and ii) decoder-only models such as Large Language Model Analysis (Llama) variants [48, 13] and Generative Pre-trained Transformer (GPT) [15] variants, like GPT-4 [14]. LLM simplification models are predominantly created through SFT; however, there has been some limited exploration of prompt-based methods and reinforcement learning.
@@
Instruction tuning allows LLMs to follow natural language instructions, facilitating prompt-based approaches such as in-context learning, chain-of-thought (CoT) learning [49, 50], and self-correction [51, 52]. These methods rely on the inherent language understanding and reasoning capabilities of LLMs [49]. CoT learning enables models to articulate step-by-step reasoning, similar to human problem solving, improving response accuracy and interpretability [49, 50]. Self-correction allows models to iteratively refine their outputs, improving the precision and reliability of their responses [52]. We present novel approaches for adapting these techniques to our exploration of SimpleDC.
@@
Recently, reinforcement learning (RL) has emerged as a promising approach for text simplification [38, 53, 54]. Many studies employ a mix of simplicity metrics and linguistic features to construct their reward functions. Zhang et al. [55] introduced an RL approach, DRESS, which includes a reward function focused on simplicity, fluency, and meaning preservation. Nakamachi et al. explored the use of grammaticality, meaning preservation, and simplicity in their reward function [53]. Similarly, Yanamoto et al. introduced an RL approach based on the difference between estimated and target difficulty levels, using a novel reward calculation method [54]. Alkaldi et al presented an RL paradigm that refines reading levels in a controlled and iterative process [56]. The TESLEA RL framework integrates heuristic and data-driven reward models to develop LLM-based text simplification models [38] using three reward models – 1) readability (FKGL), 2) relevance (cosine similarity between sentence embeddings), and 3) lexical simplicity (Zipf frequency) rewards. We adapted these components to suit our objectives, refining the reward combination method to enhance the efficacy of each element. We modified the readability reward output for better RL tuning integration and discarded the Zipf frequency-based lexical simplicity reward. Instead, we introduced a novel, non-deterministic reward function that assesses sentence simplicity. This adjustment addresses two concerns: the potential of Zipf frequency to inappropriately penalize medically significant words due to its general corpus frequency basis, and the lack of extensive parallel simplification data in the medical field.
@@ 3.1.1 SimpleDC
We introduce SimpleDC, which is a meticulously curated parallel corpus, specifically designed for text simplification research within the medical domain. It comprises educational information for patients from three high-reputation institutions: ACS, CDC, and NCI. These organizations were chosen because of their expertise and comprehensive coverage of health topics, including cancer. There are many types of cancer, each with different characteristics. The cancer educational materials on the ACS, CDC, and NCI websites exceed our annotation budget, so we focused on the subset of cancers associated with the digestive system to create a corpus that spans many cancer types that share similar terminology and topics (e.g. anatomy or symptoms). The focus on digestive cancers is also motivated by the prevalence and mortality of these cancers and the role that prevention and screening play in the outcomes [57]. SimpleDC responds to the need to make medical information more accessible and comprehensible to a broad audience, especially minoritized populations.
@@
We collected text from the ACS, CDC, and NCI web pages containing patient educational content for eight digestive cancers: anal, bile duct, colorectal, gastrointestinal, liver, pancreatic, small intestine, and stomach. For these cancer types, we collected the text from web pages that present the most introductory cancer information (e.g. About, Prevention, Screening). The collected text was cleaned and processed, including the removal of extraneous white space, adjusting punctuation, removing rich-text formatting, and parsing into sentences. We developed annotation guidelines for the simplification task and trained a team of subject matter experts (SMEs) with expertise in both cancer and patient education, including two nurse oncologists and a nurse practitioner. For each web page, we created a two-column document, where the left and right sides contained the original text with each sentence on a separate line. The annotators modified the sentences in the right column based on the annotation guidelines, to create the simplified versions. This approach allowed the annotators to focus on sentence-level simplification while considering the entire web page for context. Table 1 presents annotation examples for each source, together with the Flesch-Kincaid Grade Level (FKGL) scores for each example [58]. The FKGL is a rule-based readability formula that assesses text complexity based on the average sentence length and syllable count per word, yielding a score indicative of the US school grade level required for comprehension.
@@
In the initial training round, two annotators independently generated simplifications for each web page. To assess annotation quality, a third annotator reviewed the simplifications relative to the original text and indicated their preferred simplification as: A is better; B is better; Both are good; or None is good (assignment of annotator to I and B shuffled). In the evaluation phase, the distribution of the preferences of the third annotator was as follows: the first annotator preferred for 34% of samples, the second annotator preferred for 32% of samples, both annotators considered effective for 26% of samples, and neither simplification deemed adequate for 8% of samples. SimpleDC comprises 29 annotated web pages (14 train, 5 development, and 10 test), including 1,183 annotated sentence pairs (361 train, 294 development, and 528 test). The development and test sets were doubly annotated by two annotators and adjudicated by the third annotator to create a robust evaluation set. The training set was singly annotated. Figure 1 presents the distribution of the FKGL reading level for the original and simplified text. The original reading levels for the 3 sources are different. The simplified gold standard sentences follow a similar distribution.
@@
We also collected an additional 1,395 web pages from ACS, CDC, and NCI related to digestive cancer that were not annotated/simplified. We randomly sampled 1,000 of these unannotated web pages for use in the RL procedures described in Section .

### v3

@@ Abstract
Objective: The reading level of health educational materials significantly influences the understandability and accessibility of the information, particularly for minoritized populations. Many patient educational resources surpass widely accepted standards for reading level and complexity. There is a critical need for high-performing text simplification models for health information to enhance dissemination and literacy. This need is particularly acute in cancer education, where effective prevention and screening education can substantially reduce morbidity and mortality.
Methods: We introduce Simplified Digestive Cancer (SimpleDC), a parallel corpus of cancer education materials tailored for health text simplification research, comprising educational content from the American Cancer Society, Centers for Disease Control and Prevention, and National Cancer Institute. The corpus includes 31 web pages with the corresponding manually simplified versions. It consists of 1,183 annotated sentence pairs (361 train, 294 development, and 528 test). Utilizing SimpleDC and the existing Med-EASi corpus, we explore Large Language Model (LLM)-based simplification methods, including fine-tuning, reinforcement learning (RL), reinforcement learning with human feedback (RLHF), domain adaptation, and prompt-based approaches. Our experimentation encompasses Llama 2, Llama 3, and GPT-4. We introduce a novel RLHF reward function featuring a lightweight model adept at distinguishing between original and simplified texts when enables training on unlabeled data.
Results: Fine-tuned Llama models demonstrated high performance across various metrics. Our RLHF reward function outperformed existing RL text simplification reward functions. The results underscore that RL/RLHF can achieve performance comparable to fine-tuning and improve the performance of fine-tuned models. Additionally, these methods effectively adapt out-of-domain text simplification models to a target domain. The best-performing RL-enhanced Llama models outperformed GPT-4 in both automatic metrics and manual evaluation by subject matter experts.
@@ 1 Introduction
Accessibility and comprehensibility of patient medical educational materials are critical to improving health literacy and patient health [1]. Patient educational materials, including cancer-related materials, are essential to provide patients with the information necessary to make informed healthcare decisions. Health literacy is strongly influenced by the clarity and understandability of such materials and plays a key role in health outcomes [2]. It affects various aspects of patient health, including reducing risk factors, encouraging timely screening, and facilitating proactive health management. To ensure comprehensibility for a broad audience, the American Medical Association (AMA) and the National Institutes of Health (NIH) recommend a reading level of sixth grade or lower for patient educational materials [3, 4]. Despite the critical role of these educational materials in patient health, there remains a substantial discrepancy between the recommended and actual readability observed in practice [5]. Health educational materials from prominent sources are often at high school or college reading levels [5]. This reading complexity poses a barrier, particularly for people with lower literacy or limited health knowledge, preventing them from fully understanding and taking action on essential health information. The accessibility and comprehensibility of educational materials are particularly concerning for disseminating cancer information, where clear and understandable guidance impacts patient education and decision-making [6, 7]. Our focus on text-based cancer educational materials stems from their pivotal role in patient education and the pressing need to improve their accessibility to diverse populations.
@@
Advancements in Natural Language Processing (NLP) and Large Language Models (LLMs) offer a promising avenue to reconcile the disparity in reading levels in existing medical education materials through automatic simplification, to improve public health literacy [8]. LLMs have demonstrated proficiency in reducing linguistic complexity and preserving medical meaning [9, 10]. Finding the optimum balance of simplification and information retention is challenging, especially when the stakes of miscommunication are high [11]. Developing and evaluating data-driven text simplification models requires parallel corpora of original and simplified text. Creating these parallel corpora is extremely labor-intensive and costly. In the health domain, these challenges are exacerbated by the need for domain-specific medical expertise. To address these challenges, robust methods are needed to create high-performing simplification models with minimal training data.
@@
This paper makes significant contributions to the simplification of health information and text simplification more broadly. We focus on patient education materials related to digestive cancer, given the prevalence and mortality of these cancers [12, 13]. We introduce Simplified Digestive Cancer (SimpleDC), a novel text simplification corpus comprising patient educational materials from three prominent sources: National Cancer Institute (NCI), American Cancer Society (ACS), and Centers for Disease Control and Prevention (CDC). Although prior research has explored the simplification of health information, we are unaware of prior research focused on patient educational materials. SimpleDC is unique in its concentration on digestive cancers and includes parallel sentences (original and simplified versions), where simplifications were generated by a team of nurse oncologists and a nurse practitioner with medical and patient education backgrounds. We explore supervised fine-tuning (SFT) and reinforcement learning (RL) approaches using Llama 2 [14] and Llama 3 [15] and introduce a novel approach for reinforcement learning from human feedback (RLHF). We also explore prompt-based methods using GPT-4 [16, 17] and introduce a novel self-correction strategy that utilizes emergent reasoning capabilities. Our evaluation includes standard automatic metrics (e.g. SARI, BLEU, BERTScore, and FKGL) and a human comparison of LLM and human-generated simplifications. Automatic and manual evaluations demonstrate that the proposed RLHF reward function outperforms existing text simplification reward functions, complements SFT, and facilitates domain adaptation. Our findings advance the simplification of health text and enable the automatic creation of more accessible health information, ultimately contributing to improved health literacy. We provide the new SimpleDC corpus, trained models, and code to the research community. [^1] Link to dataset: https://github.com/mushfiqur11/simpledc-dataset.git  [^2] Link to code: https://github.com/mushfiqur11/healthLiteracy.git
@@
Problem Existing patient education materials are frequently presented at a high reading level, reducing accessibility for individuals with lower literacy.
What is already known Existing health text simplification corpora are limited, and there are knowledge gaps for low-resource text simplification related to RL, RLHF, and LLMs.
What this paper adds We present a novel text simplification corpus, SimpleDC, and introduce an innovative RLHF reward function for training LLM simplification models. The findings demonstrate the RLHF approach surpasses existing methods in performance, complements supervised fine-tuning, and supports domain adaptation.
@@ 2 Related Work
Text simplification involves the transformation of texts to improve understandability and accessibility to broad audiences. This section presents the scope of text simplification, health text simplification corpora, text simplification methods, and related RL research.
2.1 Scope of Text Simplification
Text simplification has been explored in many domains (news articles, literature, scientific articles, etc.) where complex ideas must be communicated clearly and concisely [18, 19, 20, 21, 22]. This research spans multiple languages, addressing the unique challenges and strategies needed for multilingual simplification to ensure clarity for linguistically diverse audiences [23, 24, 25]. Text simplification can be subdivided into three tasks: lexical simplification, sentence-level simplification, and paragraph/document-level simplification [26, 27, 28]. Lexical simplification commonly involves substituting words or phrases, including medical terminology [29, 30, 31, 32, 33]. However, word- or phrase-level editing limits simplification scope and cannot accommodate more comprehensive rephrasing or restructuring. Sentence- and document-level simplification enables more comprehensive text restructuring to enhance overall clarity and accessibility and typically requires parallel text corpora (complex and simplified text) for model training and evaluation [34].
2.2 Health Text Simplification
@@
Health text simplification is challenging due to medical jargon, complex sentence structures, and lengthy explanations. It has been explored through several tasks, such as generating patient-friendly summaries of clinical notes [35, 36, 37], making biomedical literature accessible to laypeople, and simplifying educational materials for patients [38, 39, 8, 40, 41, 42]. Text simplification can focus on improving readability or making stylistic modifications; however, simplification of health information requires rigorous attention to meaning preservation and medical accuracy, given the critical implications of miscommunication [8, 32]. This challenge is amplified by the diverse needs of patients, which encompasses varying levels of literacy, language background, and other factors that require a nuanced approach to achieve clarity and accuracy.
@@
Due to the labor-intensive nature of annotation, there are limited health-focused text simplification corpora [43, 44, 9, 40]. The Medical dataset for Elaborative and Abstractive Simplification (Med-EASi) dataset comprises medical text from SIMPWIKI that was simplified by a team of medical experts and layperson crowd workers at the sentence-level [43]. The Autocomplete for Medical Text Simplification (AutoMeTS) dataset consists of automatically aligned sentences from English Wikipedia and Simple English Wikipedia for medical topics [40]. Devaraj et al. provided a manually annotated parallel corpus of biomedical literature with simplified text created by domain experts [9]. We introduce a new health text simplification corpus, SimpleDC, which is based on educational patient materials from prominent information sources (ACS, CDC, and NCI) and explicitly focuses on the cancer domain.
2.3 Text Simplification Approaches
@@
Early text simplification work used heuristics and traditional machine learning techniques, including pattern-based paraphrasing [45, 46]. Contemporary approaches conceptualize text simplification similar to machine translation, employing neural language models to transform sentences into simpler counterparts [8, 47]. Document-level simplification adopts a similar machine translation approach [48, 49]. Essentially all state-of-the-art text simplification models [50, 51] are based on LLMs, including: i) encoder-decoder architectures, like Text-to-Text Transfer Transformer (T5) [52] and Fine-tuned LAnguage Net on T5 (FlanT5) [53] and ii) decoder-only models such as Large Language Model Analysis (Llama) variants [54, 14, 15] and Generative Pre-trained Transformer (GPT) [17] variants, like GPT-4 [16].
@@
LLM simplification models are predominantly created through SFT; however, there has been some limited exploration of prompt-based methods and reinforcement learning. Instruction tuning allows LLMs to follow natural language instructions, facilitating prompt-based approaches such as in-context learning, chain-of-thought (CoT) learning [55, 56], and self-correction [57, 58]. These methods rely on the inherent language understanding and reasoning capabilities of LLMs [55, 59, 60]. CoT learning enables models to articulate step-by-step reasoning, similar to human problem solving, improving response accuracy and interpretability [55, 56]. Self-correction allows models to iteratively refine their outputs, improving the precision and reliability of their responses [58]. We present novel approaches for adapting these techniques to our text simplification task.
2.4 Reinforcement Learning for Simplification
@@
RL has emerged as a promising approach for text simplification [44, 61, 62]. Many studies employ a mix of simplicity metrics and linguistic features to construct their reward functions. Zhang et al. [63] introduced an RL approach, DRESS, which includes a reward function focused on simplicity, fluency, and meaning preservation. Nakamachi et al. explored the use of rewards related to grammaticality, meaning preservation, and simplicity [61]. Similarly, Yanamoto et al. introduced an RL approach based on the difference between estimated and target difficulty levels, using a novel reward calculation method [62]. Alkaldi et al presented an RL paradigm that refines reading levels in a controlled and iterative process [64]. Luo et al demonstrated the use of controllability in the biomedical sector [65]. The TESLEA RL framework integrates heuristic and data-driven reward models to develop LLM-based text simplification models [44] using three rewards: 1) readability (FKGL), 2) relevance (cosine similarity between sentence embeddings), and 3) lexical simplicity (Zipf frequency). In our baseline experimentation, we implemented a TESLEA-inspired approach that includes the readability and relevance rewards but omits the Zipf frequency-based lexical simplicity reward to avoid penalizing medically significant but infrequent words. We introduce a new RL approach that combines the heuristic readability reward (FKGL) with a data-driven RLHF reward for assessing simplification.
@@ 3.1.1 SimpleDC
We introduce SimpleDC, which is a meticulously curated parallel corpus, specifically designed for text simplification research within the medical domain. It comprises patient educational information from three high-reputation institutions: ACS, CDC, and NCI. These organizations were chosen because of their expertise and comprehensive coverage of health topics, including cancer. There are many types of cancer, each with different characteristics. The available cancer educational materials on the ACS, CDC, and NCI websites exceed our annotation budget, so we focused on the subset of cancers associated with the digestive system to create a corpus that spans many cancer types that share similar terminology and topics (e.g. anatomy or symptoms). The focus on digestive cancers is also motivated by the prevalence and mortality of these cancers and the role that prevention and screening play in the outcomes [66]. SimpleDC responds to the need to make medical information more accessible and comprehensible to a broad audience, especially minoritized populations.
@@
We collected text from the ACS, CDC, and NCI web pages containing patient educational content for eight digestive cancers: anal, bile duct, colorectal, gastrointestinal, liver, pancreatic, small intestine, and stomach. For these cancer types, we collected the text from web pages presenting introductory cancer information (e.g. About, Prevention, Screening). The collected text was cleaned and processed, including the removal of extraneous white space, adjusting punctuation, removing rich-text formatting, and parsing into sentences. We developed annotation guidelines for the simplification task and trained a team of subject matter experts with expertise in both cancer and patient education, including two nurse oncologists and a nurse practitioner. For each web page, we created a two-column document, where the left and right sides contained the original text with each sentence on a separate line. The annotators modified the sentences in the right column based on the annotation guidelines to create the simplified versions. This approach allowed the annotators to focus on sentence-level simplification while considering the entire web page for context. Table 1 presents annotation examples for each source, together with the Flesch-Kincaid Grade Level (FKGL) scores for each example [67]. FKGL is a rule-based readability formula that assesses text complexity based on the average sentence length and syllable count per word, yielding scores indicative of the US school grade levels required for comprehension.
Original FKGL Simplified FKGL
Num. samples Num. of webpages Mean Median Mean Median Annotation type
Train 361 14 9.6 9.1 7.8 8.0 Single
Val 294 7 7.7 7.2 7.1 7.0 Double
Test 528 10 8.1 7.6 7.4 7.5 Double
Overall 1183 31 8.5 8 7.4 7.6 -
Table 2: Statistics of the proposed Simple DC dataset
@@
Since text simplification is a generative task without categorical labels, traditional inter-annotator agreement (IAA) metrics, like Cohen’s Kappa, are not well suited. Instead, we evaluated annotation quality by having two annotators create the simplified text and then having a third annotator assess their preference for simplifications. In the initial training round, two annotators independently generated simplifications for each web page. The generated texts from the two annotators were marked A and B randomly. To assess annotation quality, a third annotator reviewed the simplifications relative to the original text and indicated their preferred simplification as one of four choices – A is better; B is better; Both are good; or None is good. The distribution of the third annotator’s preferences was as follows: the first annotator was preferred for 34% of samples, the second annotator was preferred for 32% of samples, both annotators were considered effective for 26% of samples, and neither simplification was deemed adequate for 8% of samples. SimpleDC comprises 31 annotated web pages (14 train, 7 development, and 10 test), including 1,183 annotated sentence pairs (361 train, 294 development, and 528 test). The development and test sets were doubly annotated by two annotators and adjudicated by the third annotator to create a robust evaluation set. The training set was singly annotated. Figure 2 gives a detailed overview of the SimpleDC dataset. Figure 1 presents the distribution of the FKGL reading level for the original and simplified text for different data sources. The median reading level varies 7.0-9.2 across sources, and the reduction in FKGL from simplification ranged 0.6-1.0.
@@
We collected an additional 1,395 web pages from ACS, CDC, and NCI related to digestive cancer that were not annotated/simplified. We randomly sampled 1,000 of these unannotated web pages for use in the RL procedures described in Section .